### v2

@@ 4.2 Operational Model
However, mitigation does need to be perfect. It could be possible to reduce the risk of corruption when not opening all volumes by using some form of error-correction on the unopened volumes (hence sacrificing some space), and then trying to recover the volume if corruption happens. We discuss this in Section 6.4.
@@ 4.3 Security
In this section, we prove that Shufflecake achieves single-snapshot security, as defined in Section 2.3.3.
Theorem 4.1 (Single-snapshot security of Shufflecake)
The Shufflecake scheme as described in Section 4.1 is a single-snapshot (SS) secure PD scheme according to Definition 2.2.
@@ 5 Implementation and Benchmarks
We implemented the Shufflecake scheme in the C language as an open-source device-mapper-based driver for the Linux kernel. We published our code under the GPLv2+ license. The current release is v0.4.3 [45]. This section describes the programming environment and the structure of our implementation, and presents concrete performance measurements taking other popular disk encryption solutions as a baseline for comparison.
@@ 5.2 Space Utilisation
For our implementation, we fixed the block size $S_{\mathsf{B}}$ to 4096 bytes (32768 bits), so as to better amortise the per-block space overhead determined by the IVs. We chose $S_{\mathsf{L}}=256$, and $N_{\mathsf{MAX}}=15$. Since we use AES-CTR-256 as the underlying encryption scheme, we need 16-byte IVs. This led to a choice of $\Delta_{\mathsf{S}}=1$, and hence $S_{\mathsf{P}}=257$: a single 4-KiB IV block (containing 256 IVs) encrypts a 1-MiB slice. To provide a numerical summary of the space utilisation, we observe that in the case of a 1 TiB disk, the resulting theoretical maximum utilisable space is 1019.91 GiB, equal to more than 99.6% of the physical storage space.
@@ 5.3 Benchmarks
We tested our implementation looking at I/O performance and space efficiency. The test environment was a fresh installation of Ubuntu 23.04 running kernel v6.2.0 on a laptop equipped with an AMD Ryzen 7 PRO 6850U CPU with Spectre mitigations enabled, 32 GiB 4-channel 6400 MHz DDR5 RAM, and a low-power 1 TiB NVMe Micron MTFDKCD1T0TFK SSD. We tested the amount of slice fragmentation of Shufflecake (v0.4.1), its I/O performance, as well as the I/O performance of other two relevant disk-encryption tools for comparison: dm-crypt/LUKS (v6.2.0-26) and VeraCrypt (v1.25.9). All the tests were performed sequentially, on a physical primary SSD partition of size 8 GB, using the ext4 filesystem (which is the one most relevant for Shufflecake’s envisioned final use case). In the case of Shufflecake, we $\mathtt{init}$iated the partition with two volumes (one decoy and one hidden), and performed all tests on the hidden one. Analogously, in the case of VeraCrypt, we formatted the partition as a standard no-FS VeraCrypt volume, and created a 6.5 GB ext4 volume therein. In order to aid reproducibility, we also included in our implementation a suite of benchmark scripts performing the tests described here.
@@ 5.3.1 Fragmentation
For testing the I/O performance of Shufflecake against dm-crypt/LUKS and VeraCrypt, we used the fio benchmarking tool, which can flexibly measure various metrics. For each of the three disk encryption tools, we performed both random and sequential read/write operations with large amount of data on the filesystem. We fixed the same parameters for all tests, such as a queue of 32 operations and a block size of 4 kiB, which are commonly recommended to evaluate real-world performance of disks. Under these conditions, we found no observable difference in metrics between IOPs (I/O operations per second) and bandwidth (expressed in MB/s), hence we report the results looking at the bandwidth only.
@@ 6.1 Crash Consistency
As it is right now, the main obstacle to reach maturity and adoption for daily use is that Shufflecake is not crash-consistent. This means that if the program crashes during operation with one or more open volumes, data corruption is possible, because some volume state changes happen in-RAM and are cached for some time before being written on disk. This is a problem for Shufflecake, because if a crash occurs between a write of encrypted data to disk and its associated IV, the encryption becomes unrecoverable. The situation is different for solutions like LUKS or TrueCrypt, which use the XTS mode of operation and are therefore immune to this problem because they do not use explicit random per-block IVs. To fix this while keeping the property of block re-randomisation, we should make the individual logical write requests atomic, i.e. we should mask the fact that they map to several physical requests (which need to be assumed atomic): if a crash happens at any point between two of these physical requests, the old content of the logical block should still be recoverable, the disk should not be left in a “limbo” state that does not correspond to any of the logical contents written by the upper layer. We discuss here some ideas for future improvements to address these concerns.
Shufflecake incurs crash-inconsistencies when a crash happens in the time window between the update of a data block and the update of the corresponding IV block. As was discussed, Shufflecake adopts a write-on-flush approach for the IV cache, whereas data blocks are immediately written to disk, encrypted with the new IV (which is not immediately persisted on-disk); therefore, the disk is in a “vulnerable” (inconsistency-prone) state whenever the upper layer has written on a file and has not yet synced it: this is, reasonably, a large fraction of the total operating time. To solve this, it would be necessary to make the IV cache write-through; the performance impact of such a choice has not been evaluated but it would probably be heavy. It would not be sufficient, anyway, because it would only reduce the “vulnerability window” between the update of a data block and that of the IV block, it would not eliminate it.
@@ 6.2 Multi-Snapshot Security
The way it has been presented so far, Shufflecake is completely vulnerable to multi-snapshot attacks in the exact same way as TrueCrypt (and its successor, VeraCrypt): when the adversary sees “empty” slices change across snapshots, the only possible explanation is that there still is a hidden volume whose password has not been provided. We have already argued in Section 1.2 why, in practice, this level of security might be enough in most cases, and also why we believe current WoORAM-based solutions base their promise of stronger security on somewhat hard to justify assumptions. We already know [10] that achieving “complete” multi-snapshot security requires the use of WoORAMs, which have serious performance drawbacks. Regardless, as mentioned in Section 4, we designed Shufflecake with the idea of being able to add features that might help to reach an unproven, “operational” level of multi-snapshot security. In this section, we explore the possibility to achieve this goal through some separate, “orthogonal” pattern obfuscation procedure that operates independently of the main scheme and does not interfere with its single-snapshot security. We present three high-level ideas to accomplish this. They are not currently implemented, nor are they precisely specified from a conceptual point of view; instead, they are left as pointers for future research, since this area is greatly under-studied.
@@ 6.3 Shufflecake “Lite”
As we have seen in Section 6.1, crash inconsistency is a serious problem of the current embodiment of Shufflecake, caused by the use of the CTR mode of encryption which, in turn, is necessary to achieve block re-randomisation. And we have seen in Section 6.2 how this block re-randomisation is an essential ingredient for achieving some form of multi-snapshot security. However, currently no multi-snapshot security measure is implemented in Shufflecake, or even formally defined. We have already seen in Section 5.3 how the current CTR design brings some (minimal) performance hit, and how this performance hit will likely be exacerbated both by the discussed mechanisms for achieving crash consistency and by the proposed ideas for multi-snapshot security. Furthermore, we have argued in Section 1.2 how single-snapshot security might already offer a good enough security margin in many scenarios.
@@
The idea would be eventually to provide both modes for Shufflecake, “lite” and “full” (the latter including both crash consistency countermeasures and multi-snapshot capabilities), and let the user choose which one is desired during the $\mathtt{init}$ operation. In practice, one of the two would be the default choice and the other one could be selected as optional, but deciding which of the two shall be the default and which one optional will require a careful security and usability study. We leave this decision for future work, after proper confrontation with representatives of the envisioned final user demographics.
@@ 6.4.2 Corruption mitigation API
. Another possible improvement in integrating corruption-mitigation capabilities in Shufflecake is to provide an interface to communicate to upper layers (e.g., the OS, or third-party error-correction tools) whether a volume corruption happened, and possibly where. This can be done, e.g., by publishing a virtual corruption file on /sys/devices/sflc/\$VOLNAME. By detecting whether this file exists, the OS or other recovery tools can know in advance when to start a recovery process without need of user intervention. Whether a volume is corrupted or not should be detected by $\mathtt{dm\mbox{-}sflc}$ during initialisation, as explained above. Then, after the conflicting slice assignment has been resolved, Shufflecake would set a persistent corrupted flag on the affected volume. Whenever a volume is $\mathtt{open}$ed and such a set flag is read, Shufflecake would create the corruption file as explained above. In case extra information on the corruption event is known, for example the LSI of affected slices (only known the first time their corruption is detected, i.e., during the disambiguation procedure explained above), this information could be published, e.g., in /sys/devices/sflc/\$VOLNAME/corruption. This might be helpful in case certain external recovery tools work better by knowing the location of errors.
@@ Appendix 0.B Key and Glossary
$\mathsf{device}$ Block device (physical partition, USB drive, file-backed loop device, SD card, etc).
@@
$N_{\mathsf{MAX}}$ Maximum number of volumes supported per device (15 in our implementation).
@@
$\mathsf{KEK}$ Key Encryption Key, derived from a password and used to encrypt a DMB cell.
@@
$\mathsf{cell}$ DMB cell. Contains (encrypted with a KEK) the VMK for the associated volume.
@@
$\mathsf{VEK}$ Volume Encryption Key, used to encrypt position map and data slices of a volume.
@@
$\mathsf{octr}$ Occupation counter for $\mathsf{prmslices}$: the first $\mathsf{octr}$ elements of $\mathsf{prmslices}$ are $\mathsf{occupied}$.

### v3

@@ 4.2 Operational Model
However, mitigation does not need to be perfect. It could be possible to reduce the risk of corruption when not opening all volumes by using some form of error-correction on the unopened volumes (hence sacrificing some space), and then trying to recover the volume if corruption happens. We discuss this in Section 6.4.
@@ 4.3 Security
In this section, we prove that (‘Legacy’) Shufflecake achieves single-snapshot security, as defined in Section 2.3.3.
Theorem 4.1 (Single-snapshot security of Shufflecake ‘Legacy’)
The Shufflecake ‘Legacy’ scheme as described in Section 4.1 is a single-snapshot (SS) secure PD scheme according to Definition 2.2.
@@ 5 Implementation and Benchmarks
We implemented the (‘Legacy’) Shufflecake scheme in the C language as an open-source device-mapper-based driver for the Linux kernel. We published our code under the GPLv2+ license. The current release is v0.4.5 [45]. This section describes the programming environment and the structure of our implementation, and presents concrete performance measurements taking other popular disk encryption solutions as a baseline for comparison.
@@ 5.2 Space Utilisation
For our ‘Legacy’ implementation, we fixed the block size ${S_{\mathsf{B}}}$ to 4096 bytes, so as to better amortise the per-block space overhead determined by the IVs. We chose $S_{\mathsf{L}}=256$, and $N_{\mathsf{MAX}}=15$. Since we use AES-CTR-256 as the underlying encryption scheme, we need 16-byte IVs. This led to a choice of $\Delta_{\mathsf{S}}=1$, and hence $S_{\mathsf{P}}=257$: a single 4-KiB IV block (containing 256 IVs) encrypts a 1-MiB slice. To provide a numerical summary of the space utilisation, we observe that in the case of a 1 TiB disk, the resulting theoretical maximum utilisable space is 1019.91 GiB, equal to more than 99.6% of the physical storage space.
@@ 5.3 Benchmarks
We tested our ‘Legacy’ implementation looking at I/O performance and space efficiency. The test environment was a fresh installation of Ubuntu 23.04 running kernel v6.2.0 on a laptop equipped with an AMD Ryzen 7 PRO 6850U CPU with Spectre mitigations enabled, 32 GiB 4-channel 6400 MHz DDR5 RAM, and a low-power 1 TiB NVMe Micron MTFDKCD1T0TFK SSD. We tested the amount of slice fragmentation of Shufflecake (v0.4.1), its I/O performance, as well as the I/O performance of other two relevant disk-encryption tools for comparison: dm-crypt/LUKS (v6.2.0-26) and VeraCrypt (v1.25.9). All the tests were performed sequentially, on a physical primary SSD partition of size 8 GB, using the ext4 filesystem (which is the one most relevant for Shufflecake’s envisioned final use case). In the case of Shufflecake, we $\mathtt{init}$iated the partition with two volumes (one decoy and one hidden), and performed all tests on the hidden one. Analogously, in the case of VeraCrypt, we formatted the partition as a standard no-FS VeraCrypt volume, and created a 6.5 GB ext4 volume therein. In order to aid reproducibility, we also included in our implementation a suite of benchmark scripts performing the tests described here.
@@ 5.3.1 Fragmentation
For testing the I/O performance of (‘Legacy’) Shufflecake against dm-crypt/LUKS and VeraCrypt, we used the fio benchmarking tool, which can flexibly measure various metrics. For each of the three disk encryption tools, we performed both random and sequential read/write operations with large amount of data on the filesystem. We fixed the same parameters for all tests, such as a queue of 32 operations and a block size of 4 kiB, which are commonly recommended to evaluate real-world performance of disks. Under these conditions, we found no observable difference in metrics between IOPs (I/O operations per second) and bandwidth (expressed in MB/s), hence we report the results looking at the bandwidth only.
@@ 6.1 Crash Consistency
As it is right now, the main obstacle to reach maturity and adoption for daily use is that (‘Legacy’) Shufflecake is not crash-consistent. This means that if the program crashes during operation with one or more open volumes, data corruption is possible, because some volume state changes happen in-RAM and are cached for some time before being written on disk. This is a problem for (‘Legacy’) Shufflecake, because if a crash occurs between a write of encrypted data to disk and its associated IV, the encryption becomes unrecoverable. The situation is different for solutions like LUKS or TrueCrypt, which use the XTS mode of operation and are therefore immune to this problem because they do not use explicit random per-block IVs. To fix this while keeping the property of block re-randomisation, we should make the individual logical write requests atomic, i.e. we should mask the fact that they map to several physical requests (which need to be assumed atomic): if a crash happens at any point between two of these physical requests, the old content of the logical block should still be recoverable, the disk should not be left in a “limbo” state that does not correspond to any of the logical contents written by the upper layer. We discuss here some ideas for future improvements to address these concerns.
Shufflecake incurs crash-inconsistencies when a crash happens in the time window between the update of a data block and the update of the corresponding IV block. As was discussed, Shufflecake ‘Legacy’ adopts a write-on-flush approach for the IV cache, whereas data blocks are immediately written to disk, encrypted with the new IV (which is not immediately persisted on-disk); therefore, the disk is in a “vulnerable” (inconsistency-prone) state whenever the upper layer has written on a file and has not yet synced it: this is, reasonably, a large fraction of the total operating time. To solve this, it would be necessary to make the IV cache write-through; the performance impact of such a choice has not been evaluated but it would probably be heavy. It would not be sufficient, anyway, because it would only reduce the “vulnerability window” between the update of a data block and that of the IV block, it would not eliminate it.
@@ 6.2 Multi-Snapshot Security
The way it has been presented so far, Shufflecake ‘Legacy’ is completely vulnerable to multi-snapshot attacks in the exact same way as TrueCrypt (and its successor, VeraCrypt): when the adversary sees “empty” slices change across snapshots, the only possible explanation is that there still is a hidden volume whose password has not been provided. We have already argued in Section 1.2 why, in practice, this level of security might be enough in most cases, and also why we believe current woORAM-based solutions base their promise of stronger security on somewhat hard to justify assumptions. We already know [10] that achieving “complete” multi-snapshot security requires the use of woORAMs, which have serious performance drawbacks. Regardless, as mentioned in Section 4, we designed Shufflecake with the idea of being able to add features that might help to reach an unproven, “operational” level of multi-snapshot security. In this section, we explore the possibility to achieve this goal through some separate, “orthogonal” pattern obfuscation procedure that operates independently of the main scheme and does not interfere with its single-snapshot security. We present three high-level ideas to accomplish this. They are not currently implemented, nor are they precisely specified from a conceptual point of view; instead, they are left as pointers for future research, since this area is greatly under-studied.
@@ 6.3 Shufflecake “Lite”
As we have seen in Section 6.1, crash inconsistency is a serious problem of the current embodiment (‘Legacy’) of Shufflecake, caused by the use of the CTR mode of encryption which, in turn, is necessary to achieve block re-randomisation. And we have seen in Section 6.2 how this block re-randomisation is an essential ingredient for achieving some form of multi-snapshot security. However, currently no multi-snapshot security measure is implemented in Shufflecake ‘Legacy’, or even formally defined. We have already seen in Section 5.3 how the current CTR design brings some (minimal) performance hit, and how this performance hit will likely be exacerbated both by the discussed mechanisms for achieving crash consistency and by the proposed ideas for multi-snapshot security. Furthermore, we have argued in Section 1.2 how single-snapshot security might already offer a good enough security margin in many scenarios.
@@
The idea would be eventually to provide both modes for Shufflecake, “Lite” and “Full” (the latter including both crash consistency and partial multi-snapshot security), and let the user choose which one is desired during the $\mathtt{init}$ operation. In practice, one of the two would be the default mode and the other one could be selected as optional (while the ‘Legacy’ mode would eventually be deprecated), but deciding which of the two shall be the default will require a careful security and usability study. We leave this decision for future work, after proper confrontation with representatives of the envisioned final user demographics.
@@ 6.4.2 Corruption mitigation API
. Another possible improvement in integrating corruption-mitigation capabilities in Shufflecake is to provide an interface to communicate to upper layers (e.g., the OS, or third-party error-correction tools) whether a volume corruption happened, and possibly where. This can be done, e.g., by publishing a virtual corruption file on /sys/devices/sflc/\$VOLNAME. By detecting whether this file exists, the OS or other recovery tools can know in advance when to start a recovery process without need of user intervention. Whether a volume is corrupted or not should be detected by $\mathtt{dm\mbox{-}sflc}$ during instantiation, as explained above. Then, after the conflicting slice assignment has been resolved, Shufflecake would set a persistent corrupted flag on the affected volume. Whenever a volume is $\mathtt{open}$ed and such a set flag is read, Shufflecake would create the corruption file as explained above. In case extra information on the corruption event is known, for example the LSI of affected slices (only known the first time their corruption is detected, i.e., during the disambiguation procedure explained above), this information could be published, e.g., in /sys/devices/sflc/\$VOLNAME/corruption. This might be helpful in case certain external recovery tools work better by knowing the location of errors.
@@ Appendix 0.A Changelog
2024-08-23
Second revision of this document. Release reference software Shufflecake v0.4.5.
  • Fix a number of typos.
  • Clarify that this document describes the ‘Legacy’ architecture of Shufflecake, hinting at future ‘Lite’ and ‘Full’ versions of the Shufflecake scheme.
  • Clarify wording in Section 1 to answer common questions.
  • Mention that TrueCrypt and VeraCrypt also support NTFS container volumes under Windows, not only FAT, albeit with serious limitations. Thanks to Long Chongyu for pointing that out.
@@ Appendix 0.B Key and Glossary
$\mathsf{device}$ Block device (partition, USB drive, file-backed loop device, SD card, etc).
@@
$N_{\mathsf{MAX}}$ Maximum number of volumes per device (15 in our implementation).
@@
${S_{\mathsf{K}}}$ Bitsize of a generic encryption key.
$\mathsf{KEK}$ Key Encryption Key, derived from password, used to encrypt a DMB cell.
@@
$\mathsf{cell}$ DMB cell. Contains (encrypted with a KEK) the VMK for a volume.
@@
$\mathsf{VEK}$ Volume Encryption Key, encrypts a volume’s position map and data slices.
@@
$\mathsf{octr}$ Occupation counter for $\mathsf{prmslices}$: the first $\mathsf{octr}$ elements are $\mathsf{occupied}$.